# HEIGHT: HEterogeneous Interaction GrapH Transformer for Robot Navigation in Crowded and Constrained Environments

### v2

@@ VIII Conclusion
In this article, we proposed HEIGHT, a novel structured graph network architecture for autonomous robot navigation in dynamic and constrained environments. Our approach takes advantage of the graphical nature and decomposability of the constrained crowd navigation problem, introducing the following two key novelties. First, we split and process the human and obstacle representations separately. This allows the robot to effectively reason about the different geometrics and dynamics of humans and obstacles, improving its ability to navigate complex environments. Second, we propose a heterogeneous st-graph to capture various types of interactions among the robot, humans, and obstacles. The decomposition of the scenario as a heterogeneous st-graph guides the design of the HEIGHT network with attention mechanisms. Attention enables the robot to reason about the relative importance of each pairwise interaction, leading to adaptive and agile robot decision-making during navigation.
Our simulation experiments show that the HEIGHT model outperforms traditional model-based methods and other learning-based methods in terms of collision avoidance and navigation efficiency. The HEIGHT model also demonstrates improved generalization in environments with varied human and obstacle densities. In real-world environments, HEIGHT is seamlessly transferred from simulation to everyday indoor navigation scenarios without additional training, showcasting its robustness and ability to overcome the sim-to-real gap.
@@
-A Simulation experiments – Baselines
  • Dynamic window approach (DWA) [1] searches for the optimal velocity that brings the robot to the goal with the maximum clearance from any obstacle. DWA is a model-based method that only considers the current state. In addition, DWA represents both humans and obstacles as groups of small circles.
  • $\textrm{A}^{*}$+ CNN [7] is a hybrid method. With a map of the environment, $\textrm{A}^{*}$ is the global planner and generates 6 waypoints in the beginning of an episode. The inputs to the robot RL policy are a 2D LiDAR point cloud, the robot state, the waypoints, and the robot’s final goal. No human detections are used and human features are included in the point cloud. The point cloud is passed through a CNN local planner, whose output is concatenated with the embedding of other inputs. In addition to the reward function in Eq. 1, the robot receives a heading direction award if it moves towards the goal while being collision-free.
  • Decentrialized structural RNN (DS-RNN) [8] is an RL-based method that represents static obstacles as groups of small circles. In network architecture, DS-RNN only contains the RH attention and the GRU.
  • DRL-VO [30] is a hybrid method with a pure pursuit algorithm as the global planner and a RL policy as the local planner. The inputs to the planners are the positions and velocities of humans encoded in an occupancy map, a 2D LiDAR point cloud, and the robot state. The inputs are passed into a CNN network without explicit interaction modeling.
  • Homogeneous graph attention network (HomoGAT) [45] is RL-based and splits human and obstacle inputs. However, HomoGAT does not differentiate between 3 types of nodes and 3 types of edges in policy network. Instead, HomoGAT uses a single self-attention network to weigh humans, the robot, and the obstacle point cloud and feed the weighted sum of all embeddings into a GRU. HomoGAT is similar to Chen et al. [6] and Liu et al. [51] but the input, output, and training algorithm are kept the same as our method for a fair comparison.
-B Simulation experiments – Training
We train all RL methods, including all baseines and ablations except DWA, for $2\times 10^{8}$ timesteps with a learning rate $5\times 10^{-5}$. The learning rate decays at a linear rate with respect to training timesteps. To accelerate and stabilize training, we run $28$ parallel environments to collect the robot’s experiences. At each policy update, 30 steps of 6 episodes are used. We train and test all methods in a commercial desktop computer with an Intel Core i9-13900K processor, 32 GB memory, and a NVIDIA RTX 4090 GPU. Training our method takes approximately 48 hours.
@@
Effectiveness of RL planning: In Table II, DWA is a model-based approach that determines robot actions based on only the current state, which leads to the worst performance. Thus, reactive policies are not sufficient to solve our problem, justifying the necessity of long-sighted planning. In contrast, $\textrm{A}^{*}$+ CNN combines planning and RL, yet relies on occupancy maps with only obstacles and not humans. As humans move, the increasingly inaccurate occupancy map reduces the optimality of waypoints, which negatively affects overall navigation. On the other hand, RL learns to maximize the expected long-term returns based on both the current and historical states of all components in the scene. Consequently, the remaining 3 RL methods, especially HomoGAT and HEIGHT, outperform $\textrm{A}^{*}$+ CNN in most metrics and in most environments. Thus, to ensure the best navigation performance, it is important to optimize the entire planning system based on the task setting instead of optimizing only a part of it.
Difficulty of scenarios: From Table II and Table III, we observe that, besides the overfitted DS-RNN, all methods show the same trend of performance change in the 5 environments: Less crowded $>$ Less constrained $>$ More constrained $>$ More crowded. Obviously, adding more humans or obstacles increases task difficulty. But interestingly, the change in the number of humans has a larger effect on the task difficulty than the change in the number of obstacles. This phenomenon shows that avoiding collisions with humans is more difficult than avoiding obstacles in nature.
@@
The distance between the starting and the goal position of the robot ranges from 6m to 11m. The pedestrians were told to react naturally to the robot based on their own preferences. In some testing episodes, other pedestrians who were unaware of our experiment also engaged with the robot.
The configuration of hardware testing is shown in Fig. 2(b). We use a TurtleBot 2i with an RPLIDAR-A3 laser scanner. We first use the ROS gmapping package to create a map of the environment. Then, we process the map to combine small obstacles and eliminate noises. To reduce sim2real gap of the LiDAR point clouds, we use artificial point clouds obtained from localization and mapping, instead of the raw point clouds from LiDAR. The reason is that the artificial point clouds are cleaner and are not obstructed by humans. When a navigation trial begins, a user inputs the robot goal position through a keyboard. To detect human positions, we use an off-the-shelf people detector [53]. We use an Intel RealSense tracking camera T265 to obtain the pose of the robot $(p_{x},p_{y},\theta)$ and the robot wheel encoder to obtain its velocity $(u_{x},u_{y})$.

### v3

@@ VIII Conclusion
In this article, we proposed HEIGHT, a novel structured graph network architecture for autonomous robot navigation in dynamic and constrained environments. Our approach takes advantage of the graphical nature and decomposability of the constrained crowd navigation problem, introducing the following two key novelties. First, we split the human and static obstacle representations, which allows the robot to effectively reason about different collision avoidance strategies to deal with both. Second, we propose a heterogeneous st-graph to capture various types of interactions among the robot, humans, and obstacles. The heterogeneous st-graph guides the design of the HEIGHT network. Attention mechanism enables the robot to reason about the relative importance of each pairwise interaction, leading to adaptive and agile robot decisions during navigation. Our simulation benchmark shows that HEIGHT outperforms model-based and learning-based methods in terms of collision avoidance and navigation time. HEIGHT also demonstrates improved generalization when human and obstacle densities vary. HEIGHT is seamlessly transferred from simulation to real-world navigation scenarios, showcasting its robustness and deployability.
@@
Acknowledgements
We thank Zhe Huang for helpful discussions and the code of the A* planner. We thank Chen Tang, Aamir Hasan, and Rutav Shah for feedback on paper drafts. We thank colleagues at University of Illinois Urbana-Champaign and at the University of Texas at Austin who participated in real-world experiments. This material is based upon work supported by the National Science Foundation (NSF) CAREER-2143435, NSF CAREER-2046955, NSF SLES IIS-2416461, Amazon, and JP Morgan. Any opinions, findings, and conclusions expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
@@ Acknowledgements
-A Simulation experiments – Baselines
  • Dynamic window approach (DWA) [fox1997dynamic] searches for the optimal velocity that brings the robot to the goal with the maximum clearance from any obstacle. DWA is a model-based method that only considers the current state. DWA represents both humans and static obstacles as groups of small circles. We re-mapped the output of DWA to match the turtlebot kinematics and our action space.
  • Optimal Reciprocal Collision Avoidance (ORCA) [van2011reciprocal] optimizes for the optimal velocity that brings the robot to the goal by assuming that agents avoid each other under the reciprocal rule. ORCA is a model-based method that only considers the current state. In addition, ORCA represents both humans as circles and static obstacles as line segments. We re-mapped the output of ORCA to match the unicycle kinematics of our robot.
  • $\textrm{A}^{*}$+ CNN [perez2021robot] is a hybrid method with sample-based planning and RL. With a map of the environment, $\textrm{A}^{*}$ is the global planner and generates 6 waypoints in the beginning of an episode. The inputs to the RL policy are a 2D LiDAR point cloud, the robot state, the waypoints, and the robot goal. No human detections are used and human features are included in the point cloud. The point cloud is passed through a CNN, whose output is concatenated with the embedding of other inputs. In addition to the reward function in Eq. 1, the robot receives a heading direction award if it moves towards the goal while being collision-free.
  • Decentrialized structural RNN (DS-RNN) [liu2020decentralized] is an RL-based method that represents static obstacles as groups of small circles. In network architecture, DS-RNN only contains the RH attention and the GRU.
  • DRL-VO [xie2023drlvo] is a hybrid method with a pure pursuit algorithm as the global planner and an RL policy as the local planner. The inputs to the planners are the positions and velocities of humans encoded in an occupancy map, a 2D LiDAR point cloud, and the robot state. The RL network is a CNN without explicit interaction modeling.
  • Homogeneous graph attention network (HomoGAT) [velickovic2018graph] is RL-based and splits human and obstacle inputs. However, HomoGAT does not differentiate between 3 types of nodes and 3 types of edges in policy network. Instead, HomoGAT uses a single self-attention network to weigh humans, the robot, and the obstacle point cloud and feed the weighted sum of all embeddings into a GRU. HomoGAT is similar to Chen et al. [chen2019relational] and Liu et al. [liu2024sample] but the input, output, and training algorithm are kept the same as our method for a fair comparison.
Figure 9: Simulation environments for training sim2real policies in Sec. VI.
@@
Figure 12: A testing episode of our method in the real Hallway environment. The blue arrow denotes the robot path that results from its actions. The red star denotes the goal position. The turtlebot avoids two pedestrians, one after another in a narrow corridor, enters a narrow doorway, and arrives at the goal.
-B Simulation experiment setup
-B1 Training details
We train all RL methods, including all baseines and ablations except DWA, for $2\times 10^{8}$ timesteps with a learning rate $5\times 10^{-5}$. The learning rate decays at a linear rate with respect to training timesteps. To accelerate and stabilize training, we run $28$ parallel environments to collect the robot’s experiences. At each policy update, 30 steps of 6 episodes are used.
-B2 Resource usage
In simulation experiments, we train and test all methods in a commercial desktop computer with a Ubuntu 20.04 OS, an Intel Core i9-13900K processor, 32 GB RAM, and a NVIDIA RTX 4090 GPU. We report the training and testing time in simulation with the above computer in Table V. For training time, all RL methods except DRL-VO are similarly fast, while DRL-VO is the slowest due to its high-dimensional occupancy map input and the larger 2D CNN network. For inference time, HEIGHT is the slowest because the two attention networks for spatial reasoning and the GRU for temporal reasoning increases computation cost. Thus, we observe a tradeoff between performance and time efficiency. In future work, the inference latency can be reduced by (1) combining the state of multiple humans in the same group and (2) deploying transformer acceleration techniques such as key-value caching or sparse attention.
TABLE V: Comparison of training and inference time in the training environment. All RL methods are trained for $2\times 10^{8}$ steps. The inference time is averaged over 500 testing steps.
Method Training (hours) Inference (ms/step)
DWA [fox1997dynamic] N/A 1.5
ORCA [van2011reciprocal] N/A 0.3
$\textrm{A}^{*}$+ CNN [perez2021robot] 47.2 2.0
DS-RNN [liu2020decentralized] 49.0 1.0
DRL-VO [xie2023drlvo] 111.9 2.3
HomoGAT [velickovic2018graph] 47.7 2.2
HEIGHT (ours) 49.7 2.4
@@
Effectiveness of RL planning: In Table II, DWA and ORCA are model-based approaches that determines robot actions based on only the current state, which leads to the worst performances. Thus, reactive policies are not sufficient to solve our problem, justifying the necessity of long-sighted planning. As a step forward, $\textrm{A}^{*}$+ CNN combines planning and RL, yet the planner relies on occupancy maps without humans, reducing the optimality of planned waypoints and negatively affecting overall navigation. On the other hand, RL learns to maximize the expected long-term returns based on both the current and historical states of all components in the scene. Consequently, the remaining 3 RL methods, especially HomoGAT and HEIGHT, outperform $\textrm{A}^{*}$+ CNN in most metrics and in most environments. Thus, to ensure the best navigation performance, it is important to optimize the entire planning system based on the task instead of optimizing only a part of it.
Difficulty of scenarios: From Table II and Table III, we observe that, besides the overfitted DS-RNN, all methods show the same trend of performance change in the 5 environments: Less crowded >> Less constrained >> More constrained >> More crowded. Obviously, adding more humans or obstacles increases task difficulty. But interestingly, the change in the number of humans has a larger effect on the task difficulty than the change in the number of obstacles. This phenomenon shows that avoiding collisions with humans is more difficult than avoiding obstacles in nature.
@@
We train the sim2real policies in the simulators of a hallway (Fig. 9(b) in Appendix), a lounge (Fig. 9(c)), and an atrium (Fig. 9(a)) for $2\times 10^{8}$ timesteps with a decaying learning rate $5\times 10^{-5}$. The distance between the starting and the goal position of the robot ranges from 6m to 11m. The pedestrians were told to react naturally to the robot based on their own preferences. In some testing episodes, other pedestrians who were unaware of our experiment also engaged with the robot.
We use the following two sets of robot hardware:
  • In Hallway and Lounge environments, we use a TurtleBot 2i equipped with an RPLIDAR-A3 laser scanner, an Intel RealSense tracking camera T265, a Nvidia Jetson Xaiver board, and a remote host computer. We first use the ROS gmapping package to create a map of the environment. Then, we process the map to combine small obstacles and eliminate noises. To detect human positions from LiDAR point clouds, we use an off-the-shelf people detector, DR-SPAAM [Jia2020DRSPAAM]. Since DR-SPAAM tends to treat obstacles as false positive humans, we mitigate this issue by removing the point clouds that belong to pre-mapped static obstacles from the LiDAR readings, then feed the remaining point clouds that only contain humans to DR-SPAAM. We use the T265 camera to obtain the pose of the robot $(p_{x},p_{y},\theta)$ and the robot wheel encoder to obtain its velocity $(u_{x},u_{y})$. The host computer runs the robot policies and perception modules, while the TurtleBot only receives and sends ROS messages. The host computer and TurtleBot communicate through ROS.
  • In Atrium and Outdoor environment, we use a Clearpath Jackal robot equipped with an ZED 2i camera and an Ouster LiDAR. The system is built using ROS. The software stack is similar to the TurtleBot except the following differences: We use ZED 2i to detect the positions of humans and the pose of the robot. The pre-built map and robot localization are obtained from ENML algorithm [enml]. All computations are executed entirely onboard, utilizing an Intel i7-9700TE CPU and an NVIDIA RTX A2000 GPU.
[Uncaptioned image] Shuijing Liu received B.S. and Ph.D. degrees in Electrical and Computer Engineering from University of Illinois Urbana-Champaign in 2018 and 2024, respectively. She is currently a postdoctoral researcher in the Computer Science Department at The University of Texas at Austin. She also serves on the Artificial Intelligence advisory board of a startup named Municorn.ai. Her research interest is at the intersection of human-robot interaction and robot learning, with the goal of building a trustworthy partnership between humans and robots in the wild.
[Uncaptioned image] Haochen Xia is currently pursuing a B.S. degree in computer engineering, with a minor in mathematics, at the University of Illinois at Urbana–Champaign, Urbana, IL, USA. He is an undergraduate Research Assistant with the Human-Centered Autonomy Lab and the ECE Department, working under Prof. Katie Driggs-Campbell and Prof. David M. Nicol. His research interests include human–robot interaction, cognitive AI, multimodal perception, and learning-based motion planning.
[Uncaptioned image] Fatemeh Cheraghi Pouria received a B.S. degree in Mechanical Engineering from Sharif University of Technology, Tehran, Iran, in 2022. She is currently pursuing a Ph.D. degree in Electrical and Computer Engineering at University of Illinois Urbana-Champaign. Her current research interests include Human Trajectory Prediction and Robot navigation.
[Uncaptioned image] Kaiwen Hong received a B.E. degree in electrical engineering from Zhejiang University, Hangzhou, China, and the B.S. degree in electrical engineering from University of Illinois at Urbana-Champaign, Urbana, IL, USA, in 2021. He is currently pursuing a Ph.D. degree in electrical and computer engineering at the University of Illinois at Urbana-Champaign. His current research interests include robot learning and human-centered robotics.
[Uncaptioned image] Neeloy Chakraborty received B.S. and M.S. degrees in Computer Engineering from University of Illinois at Urbana-Champaign in 2021 and 2013, respectively. He is currently pursuing a Ph.D. degree in Electrical and Computer Engineering at University of Illinois at Urbana-Champaign. His current research interests include foundation model evaluation and field robotics.
[Uncaptioned image] Zichao Hu (Member, IEEE) received a B.S. degree in computer engineering from the University of Virginia, Virginia, USA, in 2022. He is currently pursuing a Ph.D. degree in computer science at the University of Texas at Austin, Texas, USA. His current research interests include diffusion policy for social navigation and whole-body control for humanoid robot object interaction.
[Uncaptioned image] Joydeep Biswas (Member, IEEE) received a B.Tech. degree in Engineering Physics from Indian Institute of Technology, Bombay in 2008, and the M.S. and Ph.D. degrees in Robotics from Carnegie Mellon University in 2010 and 2014, respectively. He is currently an Associate Professor in the Computer Science department at the University of Texas at Austin and a visiting professor at Nvidia. His research interests include robot perception, motion planning, control systems, AI, and deployed robot systems.
[Uncaptioned image] Katherine Driggs-Campbell (Member, IEEE) received a B.S.E degree in Electrical Engineering from Arizona State University in 2012, and the M.S. and Ph.D. degrees in Electrical Engineering and Computer Sciences from University of California, Berkeley in 2015 and 2017, respectively. She is currently an Associate Professor in the Electrical and Computer Engineering department at University of Illinois at Urbana-Champaign. Her research interest is developing tools for autonomous robotics that operate out in the real world with people, which is drawn from the fields of human-robot interaction, optimization, learning & AI, and control theory.